OpenStudy (anonymous):

What is the value of time, t, for the given values of distance,d, and rate r? D=315mi, r=45mi/h A. 7h B. 7mi C. 14174 h D. 14174 mi

OpenStudy (anonymous):

@shadow13 @gabby_2624

OpenStudy (mathstudent55):

\(rate~of~speed = \dfrac{distance}{time} \)

OpenStudy (shadow13):

its between b and c i believe

OpenStudy (mathstudent55):

\(r = \dfrac{d}{t} \) \(t = \dfrac{d}{r} \)

OpenStudy (mathstudent55):

The question is what is the time. The 4 choices have units of miles or hours. A time must be in hours, not miles, so the only possible answers are A. and C. Use the equation above to find out which one is correct.

OpenStudy (mathstudent55):

time = distance divided by rate

OpenStudy (anonymous):

@mathstudent55 so i think that the answer would be A

OpenStudy (shadow13):

its equals 7 h so yes u ar crecet

OpenStudy (anonymous):

ok thanks @shadow13

OpenStudy (shadow13):

welcome and this is how i got it Okay so you are traveling at 45 miles an hour and covering 315 miles. Let's divide 315 by 45 to see how many hours we spent. 315/45=7 You spent 7 hours

OpenStudy (shadow13):

anything else my friend

OpenStudy (mathstudent55):

Yes, the answer is 7 h.